OpenStudy (anonymous):

Find the volume of a sphere with diameter 21 cm.

OpenStudy (anonymous):

 r= radius

OpenStudy (anonymous):

this is what i did volume = (4/3)*pi*r^3 d=2r v=1/6 pi d^3 =1/6*pi*21^3 =4849.05cm^3

OpenStudy (anonymous):

you have to use dived 21 by 2 first then plug it in for r

OpenStudy (anonymous):

its r*r*r *4/3 that is the aproximite awnser mltyply it by pi to get the exact awnser

OpenStudy (anonymous):

@nicoleee23 does that help at all?

OpenStudy (anonymous):

yea thanks